The legs (base and height) of a right triangle are 3m and 4m. if the height of a similar right triangle is 2m, what is the other leg?
'I have an answer but I'm not sure of it.
I would be grateful for any assistance. Thank you.

Answer by Theo(13342)

if the triangles are similar, then their corresponding sides are proportional.

The side that is 3 on the first triangle is corresponding to the side that is 2 on the other triangle.

likewise the side that is 4 on the first triangle is correspond to the side that is x on the other triangle.

the ratio would be:

3/2 = 4/x

cross multiply to get:

3x = 8

divide both sides of this equation by 3 to get:

x = 8/3

The other leg is 8/3 meters.